OpenStudy (anonymous):

find nth term of ln(2x+3)

OpenStudy (anonymous):

log(2 x+3) = -sum_(n=1)^infinity((-2)^n (1+x)^n)/n for 2 |1+x|<1 log(2 x+3) = log(2 (1+x))-sum_(n=1)^infinity(-1/2)^n/(n (1+x)^n) for 2 |1+x|>1